OpenStudy (anonymous):

|2a +7|=a-4 absoulute value equations

OpenStudy (shamil98):

2a + 7 = -(a-4) 2a + 7 = a- 4 Solve for a.

OpenStudy (anonymous):

so do 7+4

OpenStudy (shamil98):

2a + 7 = a - 4 -7 -7 2a = a - 11 -a -a a = -11

OpenStudy (shamil98):

2a + 7 = -1(a-4) 2a + 7 = -a +4 solve for a from there, try doing the problem.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a=-11 a=-3

OpenStudy (shamil98):

2a + 7 = -a +4 3a = -3 a = -1.